OpenStudy (anonymous):

if A,B,C are interior angles of triangle ABC,show that cos^2 A/2+cos^2(B+C/2)=1 plss help me ma maths page is not available plsss help mee frnds....

OpenStudy (anonymous):

i will look at this. give me a second. And please verify this is the correct equation\[\cos^{2}(\frac{A}{2})+\cos^{2}(B+\frac{C}{2})=1\]

OpenStudy (anonymous):

its\[\cos^{2}(\frac{A}{2})+\cos^{2}(\frac{B+C}{2})=1\]

OpenStudy (anonymous):

yaaa

OpenStudy (anonymous):

juantweaver go on

OpenStudy (anonymous):

i am not sure what to do yet

OpenStudy (anonymous):

anyway plss show d method

OpenStudy (anonymous):

\[A+B+C=\pi\]\[\frac{B+C}{2}=\frac{\pi}{2}-\frac{A}{2}\]so\[\cos(\frac{B+C}{2})=\cos(\frac{\pi}{2}-\frac{A}{2})=\sin (\frac{A}{2})\]

OpenStudy (anonymous):

dis is too complicated i dnt wanna knw

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

thx for ur effort